lowest freezing point

which of the following aqueous liquids will have the lowestfreezing point?

0.5 m Ca(NO3)2
pure water
0.5 m Li3PO4
0.5 m C12H22O11(sucrose)
0.5 m NiSO4
Answer #1
     Depression in freezing point dependson the colligative property i.e it depends on the number of ions ofthe salt.
      So for the given salts
         0.5 mLi3PO4 gives four ions ondissociation ( 3 Li+ + PO4-3 ) = i
             Δ Tf = i Kf xmolality of the solution
Answer #2

Li3PO4 --> 3 Li+ + PO4^3- 1 0 0 1- a 3a a total no of moles after dissociation, i= 1+3a which is highest for Li3PO4 now delta Tf= i* Kf*m m and Kf for same solvent is same delta Tf is greatest for high value of i which is Li3PO4 

answered by: love_chemistry